Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Dave Cooney Ardnaglass Corrig Road Dun Laoghaire

Mark Dennehy, DURC Secretary

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC Range Officer Dinner  12/12/2001

A Chara,

The Annual Dublin University Rifle Club Range Officer’s Dinner function is to be held this Monday, December 17^{t}h, in Freemans of Dublin, on Westmoreland Street. You are cordially invited to attend, together with a guest.

Christmas Dinner Menu

Soup of the Day Poached Mussels, Shallot & Rosemary Cream Warm Duck & Citrus Salad Deep Fried Goats Cheese, Tomato Chutney Caesar Salad, Parmesan Shavings

Grilled Cod, Wild Mushroom Risotto, Chive Cream Pan fried Salmon, Crushed Herb Potatoes, Herb Oil Turkey & Ham, Chestnut Stuffing, Rosemary Sauce Sirloin Steak, Saute Potatoes, Garlic Butter, Fried Onions Lamb Shank, Champ, Charred Vegetables, Barley Jus Tagliatelle, Semi-Dried Tomatoes, Pesto Selection of Vegetables & Potatoes

Christmas Pudding, Brandy Sauce Layered Chocolate Torte, Creme Anglaise Lemon & Raspberry Tart, Mango Coulis Selection of Ice Cream, Berry Compote

£27.50 not including service charge An a la carté menu is also available.

P.S. The traditional DURC Invitational shoot is to be held as tradition dictates, on the first day College is open after Christmas, which is the 27th of December this year. Course of fire is one normal card and one rapid-fire card (10 shots in 90 seconds) and some light refreshments will be provided. We aim to start at 1400 and hope to see you there.

Mise le meas, Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Geoff Cooney Wilkinstown Navan Co. Meath

Mark Dennehy, DURC Secretary

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC Range Officer Dinner  12/12/2001

A Chara,

The Annual Dublin University Rifle Club Range Officer’s Dinner function is to be held this Monday, December 17^{t}h, in Freemans of Dublin, on Westmoreland Street. You are cordially invited to attend, together with a guest.

Christmas Dinner Menu

Soup of the Day Poached Mussels, Shallot & Rosemary Cream Warm Duck & Citrus Salad Deep Fried Goats Cheese, Tomato Chutney Caesar Salad, Parmesan Shavings

Grilled Cod, Wild Mushroom Risotto, Chive Cream Pan fried Salmon, Crushed Herb Potatoes, Herb Oil Turkey & Ham, Chestnut Stuffing, Rosemary Sauce Sirloin Steak, Saute Potatoes, Garlic Butter, Fried Onions Lamb Shank, Champ, Charred Vegetables, Barley Jus Tagliatelle, Semi-Dried Tomatoes, Pesto Selection of Vegetables & Potatoes

Christmas Pudding, Brandy Sauce Layered Chocolate Torte, Creme Anglaise Lemon & Raspberry Tart, Mango Coulis Selection of Ice Cream, Berry Compote

£27.50 not including service charge An a la carté menu is also available.

P.S. The traditional DURC Invitational shoot is to be held as tradition dictates, on the first day College is open after Christmas, which is the 27th of December this year. Course of fire is one normal card and one rapid-fire card (10 shots in 90 seconds) and some light refreshments will be provided. We aim to start at 1400 and hope to see you there.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Rhona Barry 36 Park Avenue Sandymount Dublin 4

Mark Dennehy, DURC Secretary

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC Range Officer Dinner  12/12/2001

A Chara,

The Annual Dublin University Rifle Club Range Officer’s Dinner function is to be held this Monday, December 17^{t}h, in Freemans of Dublin, on Westmoreland Street. You are cordially invited to attend, together with a guest.

Christmas Dinner Menu

Soup of the Day Poached Mussels, Shallot & Rosemary Cream Warm Duck & Citrus Salad Deep Fried Goats Cheese, Tomato Chutney Caesar Salad, Parmesan Shavings

Grilled Cod, Wild Mushroom Risotto, Chive Cream Pan fried Salmon, Crushed Herb Potatoes, Herb Oil Turkey & Ham, Chestnut Stuffing, Rosemary Sauce Sirloin Steak, Saute Potatoes, Garlic Butter, Fried Onions Lamb Shank, Champ, Charred Vegetables, Barley Jus Tagliatelle, Semi-Dried Tomatoes, Pesto Selection of Vegetables & Potatoes

Christmas Pudding, Brandy Sauce Layered Chocolate Torte, Creme Anglaise Lemon & Raspberry Tart, Mango Coulis Selection of Ice Cream, Berry Compote

£27.50 not including service charge An a la carté menu is also available.

P.S. The traditional DURC Invitational shoot is to be held as tradition dictates, on the first day College is open after Christmas, which is the 27th of December this year. Course of fire is one normal card and one rapid-fire card (10 shots in 90 seconds) and some light refreshments will be provided. We aim to start at 1400 and hope to see you there.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Ray and Collette Kane Bungalow 6 Casement Aerodrome Baldonnel Dublin 22

Mark Dennehy, DURC Secretary

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC Range Officer Dinner  12/12/2001

A Chara,

The Annual Dublin University Rifle Club Range Officer’s Dinner function is to be held this Monday, December 17^{t}h, in Freemans of Dublin, on Westmoreland Street. You are cordially invited to attend, together with a guest.

Christmas Dinner Menu

Soup of the Day Poached Mussels, Shallot & Rosemary Cream Warm Duck & Citrus Salad Deep Fried Goats Cheese, Tomato Chutney Caesar Salad, Parmesan Shavings

Grilled Cod, Wild Mushroom Risotto, Chive Cream Pan fried Salmon, Crushed Herb Potatoes, Herb Oil Turkey & Ham, Chestnut Stuffing, Rosemary Sauce Sirloin Steak, Saute Potatoes, Garlic Butter, Fried Onions Lamb Shank, Champ, Charred Vegetables, Barley Jus Tagliatelle, Semi-Dried Tomatoes, Pesto Selection of Vegetables & Potatoes

Christmas Pudding, Brandy Sauce Layered Chocolate Torte, Creme Anglaise Lemon & Raspberry Tart, Mango Coulis Selection of Ice Cream, Berry Compote

£27.50 not including service charge An a la carté menu is also available.

P.S. The traditional DURC Invitational shoot is to be held as tradition dictates, on the first day College is open after Christmas, which is the 27th of December this year. Course of fire is one normal card and one rapid-fire card (10 shots in 90 seconds) and some light refreshments will be provided. We aim to start at 1400 and hope to see you there.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Alan Lewis 6 Cherryvale Drive Glengormley Newtownabbey BT36 7UK

Mark Dennehy, DURC Secretary

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC Range Officer Dinner  12/12/2001

A Chara,

The Annual Dublin University Rifle Club Range Officer’s Dinner function is to be held this Monday, December 17^{t}h, in Freemans of Dublin, on Westmoreland Street. You are cordially invited to attend, together with a guest.

Christmas Dinner Menu

Soup of the Day Poached Mussels, Shallot & Rosemary Cream Warm Duck & Citrus Salad Deep Fried Goats Cheese, Tomato Chutney Caesar Salad, Parmesan Shavings

Grilled Cod, Wild Mushroom Risotto, Chive Cream Pan fried Salmon, Crushed Herb Potatoes, Herb Oil Turkey & Ham, Chestnut Stuffing, Rosemary Sauce Sirloin Steak, Saute Potatoes, Garlic Butter, Fried Onions Lamb Shank, Champ, Charred Vegetables, Barley Jus Tagliatelle, Semi-Dried Tomatoes, Pesto Selection of Vegetables & Potatoes

Christmas Pudding, Brandy Sauce Layered Chocolate Torte, Creme Anglaise Lemon & Raspberry Tart, Mango Coulis Selection of Ice Cream, Berry Compote

£27.50 not including service charge An a la carté menu is also available.

P.S. The traditional DURC Invitational shoot is to be held as tradition dictates, on the first day College is open after Christmas, which is the 27th of December this year. Course of fire is one normal card and one rapid-fire card (10 shots in 90 seconds) and some light refreshments will be provided. We aim to start at 1400 and hope to see you there.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Pat O’Brien Golden Hill Manor Kilbride Blessington Co. Wicklow

Mark Dennehy, DURC Secretary

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC Range Officer Dinner  12/12/2001

A Chara,

The Annual Dublin University Rifle Club Range Officer’s Dinner function is to be held this Monday, December 17^{t}h, in Freemans of Dublin, on Westmoreland Street. You are cordially invited to attend, together with a guest.

Christmas Dinner Menu

Soup of the Day Poached Mussels, Shallot & Rosemary Cream Warm Duck & Citrus Salad Deep Fried Goats Cheese, Tomato Chutney Caesar Salad, Parmesan Shavings

Grilled Cod, Wild Mushroom Risotto, Chive Cream Pan fried Salmon, Crushed Herb Potatoes, Herb Oil Turkey & Ham, Chestnut Stuffing, Rosemary Sauce Sirloin Steak, Saute Potatoes, Garlic Butter, Fried Onions Lamb Shank, Champ, Charred Vegetables, Barley Jus Tagliatelle, Semi-Dried Tomatoes, Pesto Selection of Vegetables & Potatoes

Christmas Pudding, Brandy Sauce Layered Chocolate Torte, Creme Anglaise Lemon & Raspberry Tart, Mango Coulis Selection of Ice Cream, Berry Compote

£27.50 not including service charge An a la carté menu is also available.

P.S. The traditional DURC Invitational shoot is to be held as tradition dictates, on the first day College is open after Christmas, which is the 27th of December this year. Course of fire is one normal card and one rapid-fire card (10 shots in 90 seconds) and some light refreshments will be provided. We aim to start at 1400 and hope to see you there.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Nicola Murphy 34 Mountainview Road Ranelagh Dublin 6

Mark Dennehy, DURC Secretary

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC Range Officer Dinner  12/12/2001

A Chara,

The Annual Dublin University Rifle Club Range Officer’s Dinner function is to be held this Monday, December 17^{t}h, in Freemans of Dublin, on Westmoreland Street. You are cordially invited to attend, together with a guest.

Christmas Dinner Menu

Soup of the Day Poached Mussels, Shallot & Rosemary Cream Warm Duck & Citrus Salad Deep Fried Goats Cheese, Tomato Chutney Caesar Salad, Parmesan Shavings

Grilled Cod, Wild Mushroom Risotto, Chive Cream Pan fried Salmon, Crushed Herb Potatoes, Herb Oil Turkey & Ham, Chestnut Stuffing, Rosemary Sauce Sirloin Steak, Saute Potatoes, Garlic Butter, Fried Onions Lamb Shank, Champ, Charred Vegetables, Barley Jus Tagliatelle, Semi-Dried Tomatoes, Pesto Selection of Vegetables & Potatoes

Christmas Pudding, Brandy Sauce Layered Chocolate Torte, Creme Anglaise Lemon & Raspberry Tart, Mango Coulis Selection of Ice Cream, Berry Compote

£27.50 not including service charge An a la carté menu is also available.

P.S. The traditional DURC Invitational shoot is to be held as tradition dictates, on the first day College is open after Christmas, which is the 27th of December this year. Course of fire is one normal card and one rapid-fire card (10 shots in 90 seconds) and some light refreshments will be provided. We aim to start at 1400 and hope to see you there.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Paul O’Boyle 42 Southbank Swords Co. Dublin

Mark Dennehy, DURC Secretary

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC Range Officer Dinner  12/12/2001

A Chara,

The Annual Dublin University Rifle Club Range Officer’s Dinner function is to be held this Monday, December 17^{t}h, in Freemans of Dublin, on Westmoreland Street. You are cordially invited to attend, together with a guest.

Christmas Dinner Menu

Soup of the Day Poached Mussels, Shallot & Rosemary Cream Warm Duck & Citrus Salad Deep Fried Goats Cheese, Tomato Chutney Caesar Salad, Parmesan Shavings

Grilled Cod, Wild Mushroom Risotto, Chive Cream Pan fried Salmon, Crushed Herb Potatoes, Herb Oil Turkey & Ham, Chestnut Stuffing, Rosemary Sauce Sirloin Steak, Saute Potatoes, Garlic Butter, Fried Onions Lamb Shank, Champ, Charred Vegetables, Barley Jus Tagliatelle, Semi-Dried Tomatoes, Pesto Selection of Vegetables & Potatoes

Christmas Pudding, Brandy Sauce Layered Chocolate Torte, Creme Anglaise Lemon & Raspberry Tart, Mango Coulis Selection of Ice Cream, Berry Compote

£27.50 not including service charge An a la carté menu is also available.

P.S. The traditional DURC Invitational shoot is to be held as tradition dictates, on the first day College is open after Christmas, which is the 27th of December this year. Course of fire is one normal card and one rapid-fire card (10 shots in 90 seconds) and some light refreshments will be provided. We aim to start at 1400 and hope to see you there.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Albert Jordan Bahana Farm Enniskerry Co. Wicklow

Mark Dennehy, DURC Secretary

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC Range Officer Dinner  12/12/2001

A Chara,

The Annual Dublin University Rifle Club Range Officer’s Dinner function is to be held this Monday, December 17^{t}h, in Freemans of Dublin, on Westmoreland Street. You are cordially invited to attend, together with a guest.

Christmas Dinner Menu

Soup of the Day Poached Mussels, Shallot & Rosemary Cream Warm Duck & Citrus Salad Deep Fried Goats Cheese, Tomato Chutney Caesar Salad, Parmesan Shavings

Grilled Cod, Wild Mushroom Risotto, Chive Cream Pan fried Salmon, Crushed Herb Potatoes, Herb Oil Turkey & Ham, Chestnut Stuffing, Rosemary Sauce Sirloin Steak, Saute Potatoes, Garlic Butter, Fried Onions Lamb Shank, Champ, Charred Vegetables, Barley Jus Tagliatelle, Semi-Dried Tomatoes, Pesto Selection of Vegetables & Potatoes

Christmas Pudding, Brandy Sauce Layered Chocolate Torte, Creme Anglaise Lemon & Raspberry Tart, Mango Coulis Selection of Ice Cream, Berry Compote

£27.50 not including service charge An a la carté menu is also available.

P.S. The traditional DURC Invitational shoot is to be held as tradition dictates, on the first day College is open after Christmas, which is the 27th of December this year. Course of fire is one normal card and one rapid-fire card (10 shots in 90 seconds) and some light refreshments will be provided. We aim to start at 1400 and hope to see you there.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Simon Kennedy 22 Shea’s Court Manor Street Dublin 7

Mark Dennehy, DURC Secretary

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC Range Officer Dinner  12/12/2001

A Chara,

The Annual Dublin University Rifle Club Range Officer’s Dinner function is to be held this Monday, December 17^{t}h, in Freemans of Dublin, on Westmoreland Street. You are cordially invited to attend, together with a guest.

Christmas Dinner Menu

Soup of the Day Poached Mussels, Shallot & Rosemary Cream Warm Duck & Citrus Salad Deep Fried Goats Cheese, Tomato Chutney Caesar Salad, Parmesan Shavings

Grilled Cod, Wild Mushroom Risotto, Chive Cream Pan fried Salmon, Crushed Herb Potatoes, Herb Oil Turkey & Ham, Chestnut Stuffing, Rosemary Sauce Sirloin Steak, Saute Potatoes, Garlic Butter, Fried Onions Lamb Shank, Champ, Charred Vegetables, Barley Jus Tagliatelle, Semi-Dried Tomatoes, Pesto Selection of Vegetables & Potatoes

Christmas Pudding, Brandy Sauce Layered Chocolate Torte, Creme Anglaise Lemon & Raspberry Tart, Mango Coulis Selection of Ice Cream, Berry Compote

£27.50 not including service charge An a la carté menu is also available.

P.S. The traditional DURC Invitational shoot is to be held as tradition dictates, on the first day College is open after Christmas, which is the 27th of December this year. Course of fire is one normal card and one rapid-fire card (10 shots in 90 seconds) and some light refreshments will be provided. We aim to start at 1400 and hope to see you there.

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Ian Marnane 18 Glasnevin Drive Glasnevin Dublin 11

Mark Dennehy, DURC Secretary

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: DURC Range Officer Dinner  12/12/2001

A Chara,

The Annual Dublin University Rifle Club Range Officer’s Dinner function is to be held this Monday, December 17^{t}h, in Freemans of Dublin, on Westmoreland Street. You are cordially invited to attend, together with a guest.

Christmas Dinner Menu

Soup of the Day Poached Mussels, Shallot & Rosemary Cream Warm Duck & Citrus Salad Deep Fried Goats Cheese, Tomato Chutney Caesar Salad, Parmesan Shavings

Grilled Cod, Wild Mushroom Risotto, Chive Cream Pan fried Salmon, Crushed Herb Potatoes, Herb Oil Turkey & Ham, Chestnut Stuffing, Rosemary Sauce Sirloin Steak, Saute Potatoes, Garlic Butter, Fried Onions Lamb Shank, Champ, Charred Vegetables, Barley Jus Tagliatelle, Semi-Dried Tomatoes, Pesto Selection of Vegetables & Potatoes

Christmas Pudding, Brandy Sauce Layered Chocolate Torte, Creme Anglaise Lemon & Raspberry Tart, Mango Coulis Selection of Ice Cream, Berry Compote

£27.50 not including service charge An a la carté menu is also available.

P.S. The traditional DURC Invitational shoot is to be held as tradition dictates, on the first day College is open after Christmas, which is the 27th of December this year. Course of fire is one normal card and one rapid-fire card (10 shots in 90 seconds) and some light refreshments will be provided. We aim to start at 1400 and hope to see you there.